OpenStudy (anonymous):

What is the polar form of (1, 1)?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

\(\large r=\sqrt{x^2+y^2} \) \(\large \theta=arctan\frac{y}{x} \) also, the polar form for a point is not unique.... so there are infinitely many ways to express the rectangular coordinates of (1, 1)

OpenStudy (unklerhaukus):